Solve by elimination method and what is the solution set?
3x-7y=15
-5x+4y=-25

Answer by mananth(16946)

3 x -7 y = 15 .............1
-5 x + 4 y = -25 .............2

multiply (1)by 4
Multiply (2) by 7
12 x -28 y = 60
-35 x 28 y = -175
Add the two equations
-23 x = -115
/ -23
x = 5
plug value of x in (1)
3 x + -7 y = 15
15 + -7 y = 15
-7 y = 15 -15
-7 y = 0
y = 0
CHECK

3 x + -7 y = 15
15 0 = 15
equation1 15 = 15